# TabArena: A Living Benchmark for Machine Learning on Tabular Data

### v2

@@ 3.2 Holistic Benchmarking of Peak Performance with TabArena
Abandoning the Holdout Validation Fallacy.  The rich metadata we save for TabArena allows us to study what-if cases for our evaluation design. Figure 5 shows that when using holdout validation instead of cross-validation, all models are greatly underestimated, and performance rankings are biased in favor of models that already use ensembling. This aligns with prior work [11] and demonstrates the importance of nested cross-validation in our benchmark design.
@@
Figure 5: (Left) Model Efficiency. Median times across TabPFNv2- and TabICL-compatible datasets. Training times include the time for cross-validation. (Right) Holdout vs. Cross-Validation. Predictive performance of a model with tuning and ensembling when using holdout or cross-validation.
@@
Reducing Time and Hardware Constraints for Accurate Benchmarking.  While TabArena enables a more accurate assessment of peak performance, it does so at a non-negligible computational cost. We hypothesize that prior benchmarks did not assess peak performance mainly due to time and hardware constraints. Thus, we share our result artifact to reduce time and hardware constraints for accurate benchmarking. As a result, users can cheaply compare novel models, and simulated ensembles thereof, to existing models and ensemble pipelines (see Section D.1).
@@
Furthermore, we introduce and continually maintain TabArena-Lite, a subset of TabArena that consists of all datasets with one outer fold (see Section A.4). We intend TabArena-Lite to be used in academic studies and find any novel model that outperforms other models on at least one dataset, even if it is not the state-of-the-art on average, worthy of publication.
@@
Ensembles Across Models and Individual Contributions.  While our public leaderboard summarizes model capabilities in one metric, we want to raise awareness that this is a limited perspective, wrongly suggesting that lower-ranked models are unnecessary. In Figure 6 (left), we show that a simulated ensembling pipeline using all models in TabArena outperforms all individual models and AutoGluon, further advancing the state-of-the-art. At the same time, certain individual models are more important than others for the state-of-the-art. In Figure 6 (right), we show the average weights of different models in the TabArena ensemble. Notably, models with the highest performance on the leaderboard are not necessarily the ones with the highest weights, likely because the ensemble construction favors models that perform better on the validation than test data [41, 42].
@@
A Sober Look at the “GBDT vs. Deep Learning” Debate. Prior benchmarks dedicated an enormous amount of work toward debating whether gradient-based decision trees (GBDTs) or deep learning is better across benchmarks and model studies [31, 32, 33]. We argue that the battle between GBDTs and deep learning is a false dichotomy, as both model families contribute to ensembles that strongly outperform individual model families (Figure 6). Thus, we posit that it is more important to find new models that work well in ensembles than models that beat GBDTs. TabArena enables such research by allowing users to simulate ensemble performance per and across models based on our precomputed result artifacts. Finally, to highlight, we perceive the most common reasons for practitioners to rely on GBDTs instead of deep learning to be the insufficient code quality and maintenance of deep learning methods compared to GBDT frameworks. TabArena also tackles this problem by wrapping existing deep learning methods in an easy-to-use interface, see Section D.5.
@@
Figure 6: (Left) TabArena Ensemble. Simulated performance of an ensemble using all models in TabArena compared to the leaderboard from Figure 1. (Right) Individual Contributions. Contributions of models to the TabArena ensemble in terms of average weight across all datasets.
@@ 4 Related Work
Many tabular benchmarks have been proposed over the recent years, including OpenML-CC18 [28], PMLB(Mini) [43, 44, 45], the Grinsztajn et al. [32] benchmark, TabZilla [33], the AutoML Benchmark [35], TabRepo [37], Tschalzev et al. [8], TabRed [10], the [46] benchmark, and TALENT [47]. In addition, many studies used their own, alternative evaluation frameworks [30, 20, 48, 26, 9, 5]. This shows that the community does not have a consistent understanding of how to benchmark tabular models and, more importantly, no platform to align and collectively improve benchmarking efforts. TabArena aims to provide both.
Recent studies revealed further flaws in tabular benchmarking: datasets are often outdated [7], non-IID data is used inappropriately [10], many datasets are not originally tabular [10], and inappropriate evaluation protocols are frequent [8, 11]. TabArena addresses these concerns through extensive data curation efforts and evaluation in sophisticated ensembling pipelines. As seen in Table 2, TabArena consolidates the individually outstanding aspects of previous benchmarks.
While some existing tabular benchmarks provide guidelines to add datasets or models, none include maintenance protocols or have been updated to address issues discovered in data and evaluation quality. TabArena is the first living tabular benchmark with active maintenance protocols and a public leaderboard for tabular data. Our work is inspired by benchmarking efforts from other domains, such as ChatBot Arena [39] or LiveBench [49] for LLMs and GIFT-Eval [50] for time-series forecasting.
@@ 5 Conclusion
[…]
Limitations and Societal Impact.  As we envision a living benchmarking system that will evolve over time, some limitations can be seen as future work, while others stem from fundamental trade-offs in benchmark design choices. Our (current) limitations are: (1) We use a fixed set of 200 random hyperparameter configurations to enable the study of ensemble pipelines. This prevents analyzing the variance of random hyperparameter choices [51] and studying more advanced hyperparameter optimization strategies. (2) We use a time limit per configuration; thus, our results depend on the hardware used in edge cases where the time limit is reached. Using different hardware across models (and in the future across users) reduces the comparability in such cases. (3) Our strict selection criteria for datasets makes TabArena more representative for real-world use-cases, but reduce the number of datasets drastically. We emphasize the need to work with the community to curate a more representative, high-quality collection of datasets with more diversity and statistical power. (4) Lastly, we assess predictive performance without feature engineering on top of the existing dataset state. Feature engineering could further boost predictive performance and change the model ranking.
Our work has a broader positive societal impact by improving the trustworthiness and reliability of academic benchmarks. Moreover, TabArena provides practical guidance to practitioners for small- to medium-sized IID tabular data.
@@
We thank the authors of all models with whom we were in contact for feedback and code contributions regarding the implementations of their models, namely, not among the authors of this manuscript: Yury Gorishniy; Paul Koch; Jingang Qu; and Han-Jia Ye. Furthermore, we thank the dataset donors for sharing their data, especially those under a public license. Moreover, we thank the community, particularly OpenML, Kaggle, and Hugging Face, for developing tools that facilitate the sharing of tabular data.
L.P. acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under SFB 1597 (SmallData), grant number 499552394. A.T. acknowledges funding by the German Federal Ministry of Education and Research (BMBF). F.H. acknowledges the financial support of the Hector Foundation. This research was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant number 539134284, through EFRE (FEIH_2698644) and the state of Baden-Württemberg.

### v3

@@ 3.2 Holistic Benchmarking of Peak Performance with TabArena
Abandoning the Holdout Validation Fallacy.  The rich metadata we save for TabArena allows us to study what-if cases for our evaluation design. Figure˜6 shows that when using holdout validation instead of cross-validation for model selection, all models are greatly underestimated, and performance is biased in favor of models that already use ensembling. This aligns with prior work [41, 11, 42] and demonstrates the importance of nested cross-validation in our benchmark design.
@@
Figure 6: (Left) Model Efficiency. Median training times with cross-validation across TabPFNv2- and TabICL-compatible datasets (see Figure˜4). (Right) Holdout vs. Cross-Validation. Predictive performance of a model with tuning and ensembling when using holdout or cross-validation.
@@
Reducing Time and Hardware Constraints for Accurate Benchmarking.  While TabArena enables a more accurate assessment of peak performance, it does so at a non-negligible computational cost. We hypothesize that prior benchmarks did not assess peak performance mainly due to time and hardware constraints. Thus, we share our result artifact to reduce time and hardware constraints for accurate benchmarking. As a result, users can cheaply compare novel models, and simulated ensembles thereof, to existing models and ensemble pipelines (see Section˜E.1).
@@
Furthermore, we introduce and continually maintain TabArena-Lite, a subset of TabArena that consists of all datasets with one outer fold (see Section˜A.4). We intend TabArena-Lite to be used in academic studies and find any novel model that significantly outperforms other models on at least one dataset, even if it is not the state-of-the-art on average, worthy of publication.
@@
Ensembles Across Models and Individual Contributions.  While our public leaderboard primarily ranks models, we want to raise awareness that this is a limited perspective that wrongly suggests that lower-ranked models are unnecessary. In Figure˜7 (left), we show that a simulated ensembling pipeline using all models in TabArena outperforms all individual models and AutoGluon, further advancing the state-of-the-art. In Figure˜7 (right), we show the average weights of different models in the TabArena ensemble. Notably, models with the highest performance on the leaderboard are not necessarily the ones with the highest weights, likely because the ensemble construction favors models overfitting the validation data [43, 44], such as ModernNCA and RealMLP as seen in Section˜A.6.
@@
A Sober Look at the “GBDT vs. Deep Learning” Debate. Prior benchmarks dedicated an enormous amount of work toward debating whether gradient-based decision trees (GBDTs) or deep learning is better across benchmarks and model studies [31, 32, 33]. We argue that the battle between GBDTs and deep learning is a false dichotomy, as both model families contribute to ensembles that strongly outperform individual model families (Figure˜7). Thus, we posit that it is more important to find new models that work well in ensembles than models that beat GBDTs. TabArena enables such research by allowing users to simulate ensemble performance per and across models based on our precomputed result artifacts. Finally, we hypothesize that a major reason for practitioners to rely on GBDTs instead of deep learning is the insufficient code quality and maintenance of deep learning methods compared to GBDT frameworks [45]. TabArena also tackles this problem by wrapping existing deep learning methods in an easy-to-use interface, see Section˜E.5. However, other reasons may also contribute to the slow adoption of deep learning methods, such as longer training time (especially on CPU), the lack of some functionalities, habit, the existence of more educational material for GBDTs, and overclaims of deep learning performance in academic papers.
@@
Figure 7: (Left) TabArena Ensemble. Simulated performance of an ensemble using all models in TabArena compared to the leaderboard from Figure˜1. (Right) Individual Contributions. Contributions of models to the TabArena ensemble in terms of average weight across all datasets.
@@ 4 Related Work
Many tabular benchmarks have been proposed over the recent years, including OpenML-CC18 [28], PMLB(Mini) [46, 47, 48], the Grinsztajn et al. [32] benchmark, TabZilla [33], the AutoML Benchmark [35], TabRepo [37], Tschalzev et al. [8], TabRed [10], the [49] benchmark, and TALENT [50]. In addition, many studies used their own, alternative evaluation frameworks [30, 20, 51, 26, 9, 5]. This shows that the community does not have a consistent shared understanding of how to benchmark tabular models and, more importantly, no platform to align and collectively improve benchmarking efforts. TabArena aims to provide both.
Recent studies revealed further flaws in tabular benchmarking: datasets are often outdated [7], non-IID data is used inappropriately [10], many datasets are not originally tabular [7, 10], and inappropriate evaluation protocols are frequent [8, 11]. TabArena addresses these concerns through extensive data curation efforts and evaluation in sophisticated ensembling pipelines. As seen in Table˜2, TabArena consolidates the individually outstanding aspects of previous benchmarks.
While some existing tabular benchmarks outline guidelines and commit to adding datasets or models, none include maintenance protocols or have been updated to address issues discovered in data and evaluation quality. To the best of our knowledge, the AutoML benchmark [35] and the TALENT benchmark [36] are the most similar to the vision of TabArena; both benchmarks received updates and are actively maintained. TabArena is the first living tabular benchmark for machine learning models with active maintenance protocols and a public leaderboard for tabular data.
Our work is inspired by benchmarking efforts from other domains, such as ChatBot Arena [39], RewardBench  [52], LiveBench [53], the Huggingface Open LLM Leaderboard [54, 55], and GIFT-Eval [56]. The most significant similarity between these efforts and TabArena is that we also provide a live leaderboard. Otherwise, TabArena is clearly distinct through our focus on predictive machine learning for tabular data and the resulting differences in models, datasets, and evaluation design.
@@ 5 Conclusion
We introduced TabArena, the first living benchmark for machine learning on small to medium-sized tabular data. We described the core components of TabArena: models and hyperparameter optimization, datasets, and evaluation design protocols. As part of TabArena, we share rich metadata and reproducible code with the community, allowing the community to benchmark and evaluate new models in a standardized way. We then instantiated TabArena by curating $16$ models and $51$ representative real-world tabular data tasks out of $1053$ available datasets for running a large-scale benchmark consisting of 25 $000{\sim}$25\,000\,000 individual runs that took ${\sim}15$ years of wall-clock time. To the best of our knowledge, our results are the most representative assessment of the state-of-the-art for tabular data to date. We found that deep learning models such as TabM and RealMLP, as well as foundation models for small data, perform similar to or better than gradient-boosted decision trees. Moreover, post-hoc ensembling per model and across models can dramatically improve performance.
Limitations and Societal Impact.  As we envision a living benchmarking system that will evolve over time, some limitations can be seen as future work, while others stem from fundamental trade-offs in benchmark design choices. Our (current) limitations are: (1) We use a fixed set of 200 random hyperparameter configurations to enable the study of ensemble pipelines. This prevents analyzing the variance of random hyperparameter choices [57] and studying more advanced hyperparameter optimization strategies. (2) We use a time limit per configuration; thus, our results depend on the hardware used in edge cases where the time limit is reached. Using different hardware across models (and in the future across users) reduces the comparability in such cases. (3) Our strict selection criteria for datasets makes TabArena more representative for real-world use-cases, but reduces the number of datasets drastically. We emphasize the need to work with the community to curate a more representative, high-quality collection of datasets with more diversity and statistical power. (4) Lastly, we assess predictive performance without feature engineering on top of the existing dataset state. Feature engineering could further boost predictive performance and change the model ranking.
A fundamental limitation of open-source benchmarking is that foul play or dataset contamination could compromise the leaderboard. Model providers could overfit the hyperparameters of their model on the TabArena datasets, or use them for pretraining a foundation model. We posit that keeping the TabArena alive is the best way to handle these challenges; see Section˜E.6 for a discussion.
Our work has a broader positive societal impact by improving the trustworthiness and reliability of academic benchmarks. Moreover, TabArena provides practical guidance to practitioners for small- to medium-sized IID tabular data. Lastly, while the upfront computational costs of TabArena imply a negative environmental impact, we argue that our secondary contributions, such as sharing result artifacts, will offset the negative environmental impact; we elaborate on this in Section˜D.3.
@@
We thank the authors of all models with whom we were in contact for feedback and code contributions regarding the implementations of their models, namely, not among the authors of this manuscript: Yury Gorishniy; Paul Koch; Jingang Qu; and Han-Jia Ye. Furthermore, we thank the dataset donors for sharing their data, especially those under a public license. Moreover, we thank the community, particularly OpenML, Kaggle, and Hugging Face, for developing tools that facilitate the sharing of tabular data. Finally, we thank the reviewers for their constructive feedback and contribution to improving the paper.
L.P. acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under SFB 1597 (SmallData), grant number 499552394. A.T. acknowledges funding by the Ministry of Economic Affairs, Labour and Tourism Baden-Württemberg. F.H. acknowledges the financial support of the Hector Foundation. This research was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant number 539134284, through EFRE (FEIH_2698644) and the state of Baden-Württemberg.
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction both contain our claims about contributions and scope. In detail, we claim that we (1) introduce TabArena, (2) manually curate datasets, (3) conduct a large-scale benchmark study, (4) initialize a public leaderboard, (5) assemble a team of maintainers. Moreover, we claim in the abstract and introduction that our results show: (6) the importance of validation and ensembling methods, (7) deep learning has caught up to gradient-boosted trees, (8) foundation models excel on smaller datasets, (9) ensemble across models advance state-of-the-art, and (10) that we investigate individual model contributions. Finally, we claim that (11) to share code and maintenance protocols for the living benchmark. We deliver on our claims as follows: (1,2) in Section˜2, (3,4) in Section˜3.1, (5) as part of our (visible) author list, (6-10) in Section˜3.2, and (11) in Section˜2 and Appendix˜E.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Limitations are discussed in Section˜5.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We describe our benchmark and experimental setup in Section˜2, Section˜3, provide more details in the appendix, and share all our result artifacts. Moreover, all our code is public.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We release all our data and code as part of the public TabArena framework. See Appendix˜E for user guides and links.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We describe our benchmark design and experimental setup in Section˜2, Section˜3, and provide details in the appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We include error bars in all our result plots and investigate statistical significance in Section˜A.5.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We detail information on the computer resources in Section˜3.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We believe our work conforms to the NeurIPS Code of Ethics. We were especially careful regarding the points "Copyright and Fair Use" and ethical concerns during our data curation, see Section˜B.1.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss positive societal impacts and negative societal impacts of the work in Section˜5.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We do not release new models. Moreover, we only curate existing datasets, and the code and result data artifacts we release have no perceivable risk of misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We carefully curate the assets we use and only include assets with licenses that allow our work. Moreover, we credit creators and licenses in Section˜B.2.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We detail our new dataset assets in Section˜B.2 and provide user guides for our code assets in Appendix˜E.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: LLMs are not a component of the core method in our work.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.